What should the pH of a solution of 1.0 mL of 0.1 M acetic acid and 1.0 mL of 0.1 M sodium acetate and 48.0 mL of water be? Justify your answer.
Answer #1

[CH3COONa] : 0.1 x 0.001 / 0.05 = 2x10-3 M

Ka = 1.75x10-5

pH = pKa + log [Salt]/[Acid]

pH = -log(1.75x10-5) + log(2x10-3) / (0.1)

pH = 3.05